Decide whether the ordered pair is a solution of the system:
(-6,1)
2x=-11-y
3x=-16-2y

Answer by neatmath(302)

In this case, you just need to substitute the x and y values and see if both equations are true:

2(-6)=-11-1

-12=-12 TRUE

3(-6)=-16-2(1)

-18=-16-2

-18=-18 TRUE

Thus, the point (-6,1) is a solution to this system of equations!!!

And in general, if the two lines have different slopes (which they do in this case), then this would be the ONLY solution, or the point where the two lines intersect.

I hope this helps!